# Neuro Symbolic Knowledge Reasoning for Procedural Video Question Answering with Knowledge Module Learning

### v3

@@ 3 PKG Construction and Dataset Creation
PKR-QA is a multiple-choice question answering dataset designed to evaluate reasoning over procedural knowledge in instructional videos. Each instance consists of a video segment, a question, five answer choices, and one correct answer. Unlike traditional video QA datasets that focus on grounding answers directly in the visual content of a single video, PKR-QA emphasizes reasoning over procedural knowledge that extends beyond the given video.
3.0.3 Traversal Templates for Procedural Reasoning
@@
We define questions in PKR-QA based on traversal templates which are specific reasoning patterns over PKG. For instance, the traversal $\texttt{Step}\xrightarrow{\texttt{HAS\_TOOL}}\texttt{Tool}\xrightarrow{\texttt{HAS\_PURPOSE}}\texttt{Purpose}$ corresponds to the question: “What is the purpose of the tool used in this step?”. We design 17 such traversal templates to cover a wide range of procedural reasoning types. For each template, we generate multiple question variants using GPT-4o-mini, followed by manual filtering to ensure quality and clarity. These templates form the backbone for producing diverse yet semantically consistent question-answer instances.  Figure 1 presents examples of Traversal Templates alongside their corresponding questions. The complete set of 17 question templates with example questions is provided in Table 10 (Supplementary).
3.0.4 Question and Answer Generation
@@
We generate questions by aligning each video segment with its corresponding Task and Step nodes in PKG, then apply traversal templates to form questions and retrieve answers. For each question, we sample one correct answer and four distractors. To ensure balanced answer distributions and reduce bias, we apply sampling strategies that equalize the frequency of each answer candidate across correct and incorrect options. Each question is paired with a Cypher query that retrieves supporting facts from PKG. These structured annotations serve as logical forms for KG-based evaluation, model supervision, and reasoning trace analysis.
@@
PKR-QA is designed for scenarios with limited training data but having access to structured knowledge, such as a knowledge graph. We construct a training set of 1,700 samples (100 per traversal template) and a validation set of 850 samples (50 per template). The test set contains 46,921 questions, generated from all video segments in the COIN test split. This setup supports zero- and few-shot generalization.
@@
4 Knowledge Module Learning
@@
In this section, we introduce our proposed Knowledge Module Learning (KML) for procedural knowledge reasoning, which effectively handles uncertain (probabilistic) inputs while producing interpretable reasoning outputs using a small number of trainable parameters. We train a collection of neural networks known as Neural Knowledge Module (KM) to represent each binary relation type of PKG. Then given a video and a question, we ask LLM to generate a program consisting of KM invocations to answer the questions. Then we execute those KMs sequentially with the grounded evidence extracted from the video to obtain the final answer. Next, we present the details.
Knowledge Module Learning.
@@
For each binary relation type $\mathcal{R}_{k}(\mathcal{E}_{i},\mathcal{E}_{j})$ in the PKG that maps from entity type $\mathcal{E}_{i}$ to $\mathcal{E}_{j}$, we learn a KM as a learnable neural relation ($\phi_{{\mathcal{R}}_{k}}$) with parameters ($\theta_{{\mathcal{R}}_{k}}$). KM learns to map from a given head entity $e_{i}\in\mathcal{E}_{i}$ to the corresponding set of tail entities $\mathcal{I}_{j}$ where $\mathcal{I}_{j}\subseteq\mathcal{E}_{j}$ of relation $\mathcal{R}_{k}$ as follows:
$\phi_{{R}_{k}}(~x(e_{i};~\theta_{x});~\theta_{{R}_{k}})\rightarrow\mathbf{e_{j}}$ (1)
@@
where $x(e_{i};\theta_{x})$ is a d-dimensional embeddings of the head entity and $\mathbf{e_{j}}$ is a vector that is closer to $x(e_{j})$ for all $e_{j}\in\mathcal{I}_{j}$. Here, $\mathcal{I}_{j}$ is the set of tail entities under relation $R_{k}$ for the head entity $e_{i}$ or formally $\mathcal{I}_{j}:=\left\{e_{j}\in\mathcal{E}\;\middle|\;R_{k}(e_{i},e_{j})=True\right\}$. Then the objective of KM learning is to make sure to learn $(\theta_{{R}_{k}},\theta_{x})$ such that for each head entity embedding $x(e_{i})$ can be mapped to a vector $\mathbf{e}_{j}$ that is representative of all corresponding tail entities $\mathcal{I}_{j}$. We iterate over all the triplets (including the inverse triplets) of all relation types of PKG in a batch-learning manner and train the KMs using the contrastive loss.
$log\frac{exp(\frac{\mathbf{e_{j}}\cdot x(e_{j})}{\tau})}{\sum_{\forall e_{p}\in\mathcal{B}}exp(\frac{\mathbf{e_{j}}\cdot x(e_{p})}{\tau})}$ (2)
@@
Here $\mathcal{B}$ is the set of tail entities in the batch of triplets excluding $\mathcal{I}_{j}$, $e_{j}\in\mathcal{I}_{j}$ is one of the positive target entity for $e_{i}$ under $R_{k}$, and $\tau$ is the temperature. Note that we L2 normalize all embeddings and the vector $\mathbf{e_{j}}$ before computing the contrastive loss. Contrastive loss plays a crucial role in learning neural relation functions $\phi_{{R}_{k}}$ (i.e. KMs), for modelling symbolic binary relations of the form $\mathcal{R}_{k}(\mathcal{E}_{i},\mathcal{E}_{j})$.
@@
The embedding learning function $x(;\theta_{x})$ is implemented using CLIP [39] text encoder embeddings with frozen parameters ($\theta_{x}$). Alternatively, we also learn the embedding function from scratch using standard implementations [^2] We learn these using torch.nn.Embedding. Similarly, we also learn the inverse KM for each inverse relation ${\mathcal{R}^{*}}_{k}(\mathcal{E}_{j},\mathcal{E}_{i})$ of each relation $\mathcal{R}_{k}$. Let us denote the set of all KMs by $\phi_{R}=\{\phi_{{R}_{k}}|k=1,\cdots\}$.
@@
At inference, KM takes an input embedding and maps that to an output embedding that represents a set of corresponding tail entities of that relation. For example, Table 3 shows the semantic meaning of each output embedding that maps to a set of tail entities of relation HAS_TOOL for given input embedding of the Step entity. We experimented with different neural configurations of multi-layered perceptron (MLP) and found that two-layered MLP with Tanh activation performs the best for learning KMs. Next, we present how to answer questions using KMs and LLM generated programs.
Question Answering.
@@
Given the video $v$, the question $Q$, and options $Y=\{y_{1},\cdots,y_{n}\}$ ($n=5$) we prompt a LLM $\phi()$ to find the relevant entity type ($E_{g}$) that should be grounded in the video to answer the question.
@@
where $\mathbf{x_{e}}\in\mathcal{R^{C}}$ represents the score vector for the entity type (e.g., step distribution) across all categories of that entity type. We assume there are $C$ categories for entity type $E_{g}$. To obtain estimates about the grounding entity $E_{g}$, we use ProceduralVRL (P.VRL) [64] a VLM tailored for procedural tasks.
Given the collection of Knowledge Module names ($\phi_{R}$) and the question $Q$, and the grounded entity type $E_{g}$, we use LLM (denoted by $\phi_{G}()$) to generate a list of sequence of Knowledge module invocations (known as program or $P_{1}$) to answer the question. Using a similar concept to chain-of-thought [53, 51], tree-of-thought [58], and graph-of-thought [6] we ask the LLM to generate multiple alternative programs to answer the same questions.
@@
Here $P_{l}$ is the specific module invocations and each $P_{l}$ may result in different answers following an alternative-thought of reasoning approach. The LLM module $\phi_{G}()$ can be implemented with any LLM with good reasoning capability. We use a special prompt that invokes deep understanding and consistent knowledge graph traversal and alternative thought invocations– see supplementary material for more details on the prompt. As an example, to answer the question what is an alternative tool that can be used for this step? it generates the following program of modules.
HAS_TOOL(Step) $\rightarrow$ Tool
HAS_PURPOSE(Tool) $\rightarrow$ Purpose
SIMILAR_PURPOSE(Purpose) $\rightarrow$ Purpose
PURPOSE_TO_TOOL(Purpose) $\rightarrow$ Tool
where PURPOSE_TO_TOOL(Purpose) is the inverse relation of HAS_PURPOSE(Tool). LLM may generate multiple programs that leads to the answer, typically 3-5 alternatives for complex problems. To predict the answer, we execute each program in order, inputting the grounded entity representation $z_{i}$ into the first module. For each program ($P_{1}$ to $P_{l}$), $z_{i}$ is fed into the first module to obtain intermediate embedding, invoking all modules sequentially. For example $z_{j}=\phi_{{r}_{i}}(z_{i})$ then $z_{k}=\phi_{{r}_{j}}(z_{j})$ and finally $z_{f}=\phi_{{r}_{k}}(z_{k})$. Therefore the final embedding representing the answer to the given question is $z_{f}$. We can also inspect the meaning of each intermediate representation as shown in Table 3 allowing more interpretable reasoning that can handle uncertain inputs.
@@
Inference Given the options $Y=\{y_{1},\dots,y_{n}\}$, we obtain the embeddings of $Y$, i.e., $x(Y)=[x(y_{1}),x(y_{2}),\cdots x(y_{n})]$. Then we compute the cosine similarity between the final embedding $z_{f}$ and $x(Y)$ and apply softmax to predict the answer index for each program. When there are alternative programs, we take the maximum score from all programs as the final answer.
@@
VQA Training for KML. We also fine-tune the KMs using a few examples of the video-question-answer using our dataset. We compute the cosine similarity between $z_{f}$ and $x(Y)$ and apply softmax to predict the answer index.
$\hat{y}=softmax\left(cosine\_sim\left(x(Y),z_{f}\right)\right)$ (8)

### v5

@@ IV Question Answering Dataset Creation
PKR-QA is a multiple-choice question answering dataset designed to evaluate reasoning over procedural knowledge in instructional videos. Each instance consists of a video segment, a question, five answer choices, and one correct answer. Unlike existing video QA datasets that focus on grounding answers directly in the visual content of a single video, PKR-QA emphasizes reasoning over procedural knowledge that extends beyond the given video.
IV-A Traversal Templates for Procedural Reasoning
@@
We define questions in PKR-QA based on traversal templates which are specific reasoning patterns over PKG. For instance, the traversal $\texttt{Step}\xrightarrow{\texttt{HAS\_TOOL}}\texttt{Tool}\xrightarrow{\texttt{HAS\_PURPOSE}}\texttt{Purpose}$ corresponds to the question: “What is the purpose of the tool used in this step?”. We design 17 such traversal templates to cover a wide range of procedural reasoning types. For each template, we generate multiple question variants using GPT-4o-mini, followed by manual filtering to ensure quality and clarity. These templates form the backbone for producing diverse yet semantically consistent question-answer instances.  Figure 1 presents examples of Traversal Templates alongside their corresponding questions. The complete set of 17 question templates with example questions is provided in the Table X - Appendix.
IV-B Question and Answer Generation
We generate questions by aligning each video segment with its corresponding Task and Step nodes in PKG, then apply traversal templates to form questions and retrieve answers. For each question, we sample one correct answer and four distractors. When generating answer options for each question, relying solely on random sampling can lead to biased options. For instance, frequent terms (e.g., tasks, grounded tools, or objects) are more likely to appear in the correct options compared to infrequent terms, which tend to appear only in the incorrect options. This creates a shortcut for models to identify the correct answer without engaging in true information retrieval and reasoning. To address this issue, we design tailored sampling strategies for each traversal template, ensuring a relative-balanced distribution of terms between correct and incorrect answer options.
@@ IV-B Question and Answer Generation
Additionally, each question is paired with a Cypher query that retrieves supporting facts from PKG. These structured annotations serve as logical forms for KG-based evaluation, model supervision, and reasoning trace analysis.
@@
PKR-QA is designed for scenarios with limited training data but having access to structured knowledge, such as a knowledge graph. We construct a training set of 1,700 samples (100 per traversal template) and a validation set of 850 samples (50 per template). The test set contains 46,921 questions, generated from all video segments in the COIN test split. This setup supports zero-shot and few-shot generalization.
@@
IV-C Rephrasing Question Templates
To ensure a diverse range of question phrasings in our dataset, we performed a rephrasing process. Specifically, we prompted GPT-4o-mini to generate up to 30 semantically diverse and contextually appropriate rephrasing for each traversal-based natural language question template. This step leverages the LLM’s capability to enhance linguistic variety while preserving meaning. To ensure quality and consistency, all rephrased questions were manually reviewed to confirm that the original intent was retained. The prompt is shown in in Figure 9 - Appendix.
V Knowledge Module Learning
@@
In this section, we introduce our Knowledge Module Learning (KML) for procedural knowledge reasoning, which effectively handles uncertain (probabilistic) inputs while producing interpretable reasoning outputs using a small number of trainable parameters. We train a collection of neural networks known as Neural Knowledge Module (KM) to represent each binary relation type of PKG. Then given a video and a question, we ask LLM to generate a program consisting of KM invocations to answer the questions. Then we execute those KMs sequentially with the grounded evidence extracted from the video to obtain the final answer. Figure 2 shows the overall of our Knowledge Module Learning (KML) framework. Next, we present the details.
Figure 2: Knowledge Module Learning (KML) for knowledge graph-based procedural video question answering. Given a video and a question, the framework first identifies the entity type in PKG (e.g. “step”) that should be detected in the video and then recognizes the entity type instances (e.g., step: “wrap the pipe band”) using a VLM. Based on the question and grounded entity type, the answer module generator produces the module sequence. Finally, KML executes the sequence of Knowledge Modules to answer the question. In this example, it first identifies the next step and then determines the tool required for that step, which is a “wrench”.
V-A Knowledge Module Learning
@@
For each binary relation type $\mathcal{R}_{k}(\mathcal{E}_{i},\mathcal{E}_{j})$ in the PKG that maps from entity type $\mathcal{E}_{i}$ to $\mathcal{E}_{j}$, we learn a KM as a learnable neural relation ($\phi_{{\mathcal{R}}_{k}}$) with parameters ($\theta_{{\mathcal{R}}_{k}}$). KM learns to map from a given head entity $e_{i}$ to the corresponding set of tail entities $\mathcal{I}_{j}$ where $\mathcal{I}_{j}\subseteq\mathcal{E}$ of relation $\mathcal{R}_{k}$ as follows:
$\phi_{{R}_{k}}(~h(e_{i};~\theta_{h});~\theta_{{R}_{k}})\rightarrow\mathbf{e_{j}}$ (1)
where $h(e_{i};\theta_{h})$ is a d-dimensional embedding function.
@@
The entity $e_{i}$ is mapped to $\mathbf{e_{j}}$, a vector that is closer to $x_{j}=h(e_{j})$ for all $e_{j}\in\mathcal{I}_{j}$ (the corresponding set of tail entities). Formally, $\mathcal{I}_{j}$ is the set of tail entities under relation $R_{k}$ for the head entity $e_{i}$, i.e. $\mathcal{I}_{j}:=\left\{e_{j}\in\mathcal{E}\;\middle|\;R_{k}(e_{i},e_{j})=True\right\}$.
The objective of Knowledge Module Learning (KML) is to learn a relation-specific neural mapping function $\phi_{R_{k}}$ (cf. Equation 1) such that, given a head entity $e_{i}$, the mapped representation $\phi_{R_{k}}(h(e_{i}))$ is closer (under cosine similarity) to embeddings of all valid tail entities than to embeddings of invalid entities.
Let $X=\{x_{1},x_{2},\dots,x_{n}\}\subset\mathbb{R}^{D}$ denote the set of learned entity embeddings ($x_{j}=h(e_{j},\theta_{h})$) in the knowledge graph, where each $x\in X$ is $\ell_{2}$-normalized, i.e., $\|x(e)\|_{2}=1$. For a fixed relation $R_{k}$ and a head entity embedding $x\in X$, we define $Y(x)\subseteq X$: the set of positive tail entity embeddings satisfying $(e_{i},R_{k},e_{j})$. Let $U(x):=X\setminus Y(x)$: the set of negative tail entity embeddings and $z:=\phi_{R_{k}}(x)$: the output of the relation-specific knowledge module where $\hat{z}:=z/\|z\|_{2}$ is the normalized module output. We define cosine similarity as $s(a,b):=\langle a,b\rangle,\qquad\|a\|_{2}=\|b\|_{2}=1,$ so that $s(a,b)\in[-1,1]$.
We iterate over all the triplets (including the inverse triplets) of all relation types of PKG in a batch-learning manner and train the KMs using the contrastive loss.
$\mathcal{L}(x)\;=\;-\log\frac{\sum\limits_{y\in Y(x)}\exp\!\big(\frac{s(\hat{z},y)}{\tau}\big)}{\sum\limits_{v\in X}\exp\!\big(\frac{s(\hat{z},v)}{\tau}\big)}.$ (2)
@@
Here $\tau$ is the temperature. Contrastive loss plays a crucial role in learning neural relation functions $\phi_{{R}_{k}}$ (i.e. KMs), for modelling symbolic binary relations of the form $\mathcal{R}_{k}(\mathcal{E}_{i},\mathcal{E}_{j})$ as later shown by the theoretical analysis on the learning of such mappings in Section VI-A.
@@
The embedding learning function $h(;\theta_{h})$ is implemented using CLIP [radford2021learning] text encoder embeddings with frozen parameters ($\theta_{h}$). Alternatively, we also learn the embedding function from scratch using standard implementations [^3] We learn these using torch.nn.Embedding. Similarly, we also learn the inverse KM for each inverse relation ${\mathcal{R}^{*}}_{k}(\mathcal{E}_{j},\mathcal{E}_{i})$ of each relation $\mathcal{R}_{k}$. Let us denote the set of all KMs by $\phi_{R}=\{\phi_{{R}_{k}}|k=1,\cdots\}$.
@@
At inference, KM takes an input embedding and maps that to an output embedding that represents a set of corresponding tail entities of that relation. For example, Table VI shows the semantic meaning of each output embedding that maps to a set of tail entities of relation HAS_TOOL for given input embedding of the Step entity. We experimented with different neural configurations of multi-layered perceptron (MLP) and found that two-layered MLP with Tanh activation performs the best for learning KMs. As shown later in Section VI-B, controlling the Lipschitz constant of the neural mapping function in KMs is crucial. Therefore, both the choice of activation functions and the design of the MLP are carefully considered. Next, we present how to answer questions using KMs and LLM generated programs.
V-B Question Answering suing KML
@@
Given the video $v$, the question $Q$, and options $O=\{o_{1},\cdots,o_{n}\}$ ($n=5$) we prompt a LLM $\phi()$ to find the relevant entity type ($E_{g}$) that should be grounded in the video to answer the question.
@@
where $\mathbf{x_{e}}\in\mathcal{R^{C}}$ represents the score vector for the entity type (e.g., step distribution) across all categories of that entity type. We assume there are $C$ categories for entity type $E_{g}$. To obtain estimates about the grounding entity $E_{g}$, we use ProceduralVRL (P.VRL) [zhong2023learning] a VLM tailored for procedural tasks.
Given the collection of Knowledge Module names ($\phi_{R}$) and the question $Q$, and the grounded entity type $E_{g}$, we use LLM (denoted by $\phi_{G}()$) to generate a list of sequence of Knowledge module invocations (known as program or $P_{1}$) to answer the question. Using a similar concept to chain-of-thought [wei2022chain, wang2022self], tree-of-thought [yao2023tree], and graph-of-thought [besta2024graph] we ask the LLM to generate multiple alternative programs to answer the same questions.
@@
Here $P_{l}$ is the specific module invocations and each $P_{l}$ may result in different answers following an alternative-thought of reasoning approach. The LLM module $\phi_{G}()$ can be implemented with any LLM with good reasoning capability. We use a special prompt that invokes deep understanding and consistent knowledge graph traversal and alternative thought invocations. We use LLM to model the  Equation 3 and  Equation 5 (See Figure 8 - Appendix).
As an example, to answer the question what is an alternative tool that can be used for this step? it generates the program shown in Figure 3.
Figure 3: Example of logical programs generated by LLM.
Here PURPOSE_TO_TOOL(Purpose) is the inverse relation of HAS_PURPOSE(Tool). LLM may generate multiple programs that leads to the answer, typically 3-5 alternatives for complex problems. To predict the answer, we execute each program in order, inputting the grounded entity representation $z_{i}$ into the first module. For each program ($P_{1}$ to $P_{l}$), $z_{i}$ is fed into the first module to obtain intermediate embedding, invoking all modules sequentially. For example $z_{j}=\phi_{{r}_{i}}(z_{i})$ then $z_{k}=\phi_{{r}_{j}}(z_{j})$ and finally $z_{f}=\phi_{{r}_{k}}(z_{k})$. Therefore the final embedding representing the answer to the given question is $z_{f}$. We also inspect the meaning of each intermediate representation as shown in Table VI allowing more interpretable reasoning that can handle uncertain inputs.
@@
Inference Given the options $O=\{o_{1},\dots,o_{n}\}$, we obtain the embeddings of $O$, i.e.,
$h(O)=[h(o_{1}),h(o_{2}),\cdots h(o_{n})].$ (8)
@@
Then we compute the cosine similarity between the final embedding $z_{f}$ and option embeddings $h(O)$ and apply softmax to predict the answer index for each program. When there are alternative programs, we take the maximum score from all programs as the final answer.
V-C VQA Training for KML.
@@
We also fine-tune the KMs using a few examples of the video-question-answer using our dataset. We compute the cosine similarity between $z_{f}$ and $h(O)$ and apply softmax to predict the answer scores.
$\hat{y}=softmax\left(cosine\_sim\left(h(O),z_{f}\right)\right)$ (9)